In a college, out of 4320 students, 2300 are girls. Find the ratio of
A number of boys to the total number of students.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given
Total number of students = 4320
Number of girls = 2300
Number of boys = 4320–2300
= 2020
Ratio of number of boys to the total number of students = 2020 / 4320
= (2 × 2 × 5 × 101) / (2 × 2 × 5 × 216)
= 101 / 216

Thanks